Please help,l have 1L of 96% alcohol and l want to make a drink of 30% alcohol,how many liters of water do l add?
Answer by Alan3354(69443)

Please help,l have 1L of 96% alcohol and l want to make a drink of 30% alcohol,how many liters of water do l add?
-------------
1*96 + x*0 = (1+x)*30
96 = 30x + 30
x = 66/30 = 2.2
Add 2.2 liters